OpenStudy (anonymous):

simplest form of (2^3+4^3)^-1/9

OpenStudy (anonymous):

\[(2^{3}+4^{3})^{\frac{ -1 }{ 9 }}\]

OpenStudy (anonymous):

\(2^3=8,4^3=64,8+64=72\) is a start

OpenStudy (anonymous):

I've done this over and over on my calculator and I cant get it

OpenStudy (anonymous):

\[72^{-\frac{1}{9}}=\frac{1}{\sqrt[9]{72}}\]

OpenStudy (anonymous):

if you want a decimal with a calculator, then try this http://www.wolframalpha.com/input/?i=%282^3%2B4^3%29^-1%2F9